Mulaudzi liquor operates in a competitive industry to produce alcohol at a constant marginal cost of R28 per unit. Immediately when the industry is monopolized, marginal cost rise to R35 per unit because of the R5 per unit must be paid to lobbyists to ensure that only Mulaudzi firm receive liquor license


Qd=1300-15P


And Marginal revenue curve by


MR= 15-Q/10



a) Calculate the monopoly quantity


1
Expert's answer
2021-07-07T09:00:49-0400

The monopoly firm will produce at a point where MR=MC for profit maximization

"MR=MC\\15-\frac{Q}{10}=35\\-\frac{Q}{10}=35-15\\-\frac{Q}{10}=20\\-Q=20\times 10\\-Q=200\\Q=-200"